8/Dec/99 Eve Shift

R Tschirhart (tsch@fnal.gov)
Wed, 8 Dec 1999 23:54:26 -0600 (CST)


Shift Summary for eve shift, 8/Dec/99

On Shift: Ruth & Stephen, BobT

Hours of beam available at all intensities: 0.00
Hours of beam available at nominal intensity: 0.00
Hours of E832 data (special runs not included): 0.00
Hours of special runs (laser scans, muon runs...): 8.00
========================================================================
Accounting of lost time: Magnet change

========================================================================
Runs ended this shift:
Diagnostic cosmic runs, a series of laser scans.
Current cosmic run 15165

Tapes brought to FCC this shift: None

=======================================================================
Accesses during shift: Open access.

Activities during the shift while waiting for beam:

Rick Coleman and Rick Ford add a 4th accidental counter to the
90-degree monitor telescope, which will be tested out with the
return of beam.

Ron Ray replaces the last remaining untested segment of the HV
cable plant for CsI 2893. First look at laser scan data is
encouraging; noise sources improved, possibly fixed. No
outstanding problems remain in the dynode-signal/HCC system!

DAMP DAQ monitor dies, Stephen restarts DAMP after consultation
with Margaret.

KTeV3b fills up due to ongoing Trigger-1 split. Problem was
traced to not accounting for quiescent laser data that has been
accumulating on KTeV3b for a series of long cosmic runs to
measure the crystal uniformities. Some split data had to be
deleted, recovery in progress.

========================================================================
Follow up Notes on Previous Shift


========================================================================
Items to Follow up on Next Shift

Gas Check list.